# Out of Sight, Still in Mind: Reasoning and Planning about Unobserved Objects with Video Tracking Enabled Memory Models

### v1

@@ I Introduction
For robots to assist humans in their daily lives in roles such as home assistants and caregivers for elders, they must be able to reason about objects not observed in their current perceptual data. For example, if asked to retrieve an apple stored inside a cabinet, the robot should remember where it was placed and that it must first open the cabinet to grasp the object. Further if the robot picks up a box containing objects it previously placed inside the container, it should know that these objects will move with the box, while those sitting near the box will not. We desire for robots to maintain these models over many separate task assignments from their operators in order to enable long-term autonomy.
The problem of maintaining a persistent model of a robot’s world from partial observations has a rich history in robotics, primarily in the form of mapping [1, 2, 3, 4, 5]. While some attempts have been made to incorporate object-level semantics into sensor-derived maps [6, 7, 8] these approaches primarily build immutable, monolithic models not suitable for complex manipulation planning such as multi-object rearrangement [9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. Such reasoning can enable robots to predict preconditions for actions [12], achieve language goals [13], and achieve logical goal relations [18, 15]. Thus a problem arises in how to build a consistent estimate of the world state amenable to contemporary, learning-based approaches to manipulation planning.
In contrast, the standard learning-based approach to persistent memory treats the problem as a sequence prediction task, typically using autoregressive recurrent or transformer-based neural networks [19, 20, 21, 22, 23]. While these models have shown some interesting results, we hypothesize that they will have difficulty maintaining long-term memory of objects that remain occluded through many updates to a recurrent memory. One approach to overcome this issue would be to maintain the entire history of robot observations in order to use as input to a model (e.g. transformer) that could selectively attend to the relevant bits of input to determine a coherent state estimate. This is obviously untenable for long-horizon tasks as the input would grow linearly with time, inevitably becoming too large to efficiently manage.
Figure 2: Overview of our approaches. As the robot takes action over time, some objects may disappear and reappear, and some objects may newly appear in the scene. In this paper, we propose two types of object-oriented memory, called DOOM and LOOM, that enable the robot to plan with occluded and newly appeared objects. DOOM and LOOM utilize a UVOS algorithm to keep track of the current object list and update object memory slots based on the occlusion status of each object accordingly. Best viewed in color.
@@
In this paper we advocate for using a UVOS algorithm to explicitly manage our object-oriented memory. We hypothesize that explicit encoding of objects into the memory will be more robust than implicit, autoregressive models in managing long-term history and successfully performing downstream planning. We desire that the representation provide sufficient information to enable prediction of inter-object and object-environment relations, while also having the ability to predict the effects of the robot’s actions on all objects in memory.
Specifically, we examine incorporating an explicit UVOS-based memory model into the framework from Huang et al. [17, 18], which was an effective framework to learn relational dynamics across varying object and environments. Key to its success is the ability to encode a variable number of objects for a given observation using a graph neural net [17] or transformer-based encoder [18]. Prediction of relations enables the model’s use in logic-based task planning [29], where relations have proved an effective means of communication between robots and humans [15, 13, 10, 30, 12, 17]. However, the existing framework assumes all relevant objects to be observable; and thus fails to successfully plan and execute in the scenarios shown in Fig. 1.
@@
To validate our proposed approaches, we test on varying tasks involving occluded objects, novel objects, and objects that reappear. We show some example tasks in Fig. 1. We show that both forms of memory-based state augmentation can reliably plan actions to change inter-object relations involving currently unobserved objects, such as placing object $A$ to the left of currently occluded object $B$. Furthermore we compare our explicit memory approach to a baseline implicit memory model in a scenario specifically designed to test the models’ long-term memory. Our experiments show that the implicit memory model fails to effectively reason about occluded objects, while our explicit, UVOS-based memory models continue to plan reliably. Our work thus represents the first successful approach of using memory models for reasoning about and planning to inter-object goal relations for an a priori unknown, variable number of objects. For more real-world robot executions, additional results, and supplemental material, see our website https://sites.google.com/view/rdmemory.
@@ II Related work
Reasoning about object permanence is an important capability for robot manipulation [31, 22, 32, 33]. Xu et al. [31] and Ebert et al. [22] are the first ones to propose deep learning models to reason about occluded objects and do downstream planning with the learned dynamics models. However, both of these approaches assume goal images for planning which may not always be available from human operators. They additionally examine only planar pushing tasks with only a single moving object at each time. Our work examines a much more diverse set of tasks and skills, while also requiring only logic-based goal representations. Curtis et al. [33] propose a system with modules that can estimate affordances and properties to perform multi-step manipulation tasks with unknown and occluded objects. However, they assume complete object shape and have many engineered modules including the affordance module.
@@
Other works have also investigated using graph neural networks [42, 17, 10, 43, 44], transformers [13, 45, 46, 47, 48], and diffusion models [9, 49, 50] to reason about and manipulate multiple objects. However, none have examined explicit memory management. Since multi-object manipulation usually requires a multi-step process, task and motion planning has shown promise in solving long-horizon, multi-object planning problem [51, 52, 53, 54, 18, 55]. These planners either require all objects to be known a priori in order to perform tracking for belief space planning [53] or ignore occlusion of unknown objects.
@@ III Approach
We assume the robot perceives the world as a point cloud, $Z_{t}$, at each timestep $t$. The robot then takes action $A_{t}$ and receives subsequent observation $Z_{t+1}$. At new observation $Z_{t+1}$, some objects may have become occluded and other, new objects may appear. Based on the history of observations and actions $Z_{0:t}$, $A_{0:t}$ we would like the robot to plan to achieve a goal, potentially involving previously observed, but currently occluded objects.
We define the goal as a logical conjunction of $M$ desired object and environment relations, $\mathbf{g}=r_{1}\land r_{2}\land...\land r_{M},r_{j}\in\mathcal{R}$, where $\mathbf{g}$ denotes the goal conjunction, $r_{j}$ represents each goal relation, and $\mathcal{R}$ denotes all possible relations. Our robot is given a set of $L$ parametric action primitives $\mathcal{A}=\{A_{1},\ldots,A_{L}\}$ where $A_{l}$ defines a skill, which has associated continuous skill parameters $\theta_{l}$. For example, a push skill is defined with parameters encoding the push direction and length, or a pick-and-dump skill is defined with parameters encoding the grasp pose and dump pose. The robot’s planning task is defined as finding skills and skill parameters $\tau=((A_{0},\theta_{0}),\ldots,(A_{H-1},\theta_{H-1}))$ that, when sequentially executed, transform the objects so they satisfy all desired object and environment relations in the goal $\mathbf{g}$.
@@
Segmentation and Tracking We first perform UVOS using [28] and obtain $N$ object and environment segments $O^{i}_{t}\subset Z_{t},i=1,2,...,N$ for a timestep $t$. The UVOS approach checks for new objects by examining the objectness of object proposals on key frames, when the robot finishes the execution of a skill. Then object segments that can be matched with previously tracked objects are assigned consistent IDs, while those that do not match any previous object are assigned new IDs. For non-key frames, the algorithm tracks all segments with a space-time transformer model [41] where each object has its own memory. It stores prior appearances of all objects hence can easily re-identify previously observed and disappeared objects. To familiarize the UVOS model with the robot, it is fine-tuned using annotations of the robot arm and objects from YCB-Video dataset [56]. Nonetheless, it still demonstrates generalization to novel object types, like the tall cups (Row. 1), apple and orange (Row. 2) shown in Fig. 1 that it was never trained on. For environment segments, we use RANSAC [57] to find planar surfaces.
Pose and relation detection To jointly reason about the pose of each object and the relations between each pair of objects, we first process each segment $O^{i}_{t}$ with a point cloud encoder [58] to get a feature vector per segment with $P^{i}_{t}=E_{p}(O^{i}_{t})$. We use a learned positional embedding to encode the ID of each object as $I_{i}$. We randomly generate the ID for each object during training to improve generalization [59]. We get the latent space at current step $t$ as $\mathbf{x}^{i}_{t}=E(P^{i}_{t}\oplus I_{i})$, where $E$ is a transformer encoding the interaction between different segments. To predict object pose and relations at the current step, we define three different decoders: the pose decoder $D_{p}$, the relation decoder $D_{r}$, and the environment identity decoder $D_{e}$. The third decoder classifies whether a given segment is movable or not (e.g. is a table or shelf) as $\hat{e}^{i}_{t}=D_{e}(\mathbf{x}^{i}_{t})$. The associated predictions of the other decoders are $\hat{p}^{i}_{t}=D_{p}(\mathbf{x}^{i}_{t})$ and $\hat{r}^{ij}_{t}=D_{r}(\mathbf{x}^{i}_{t},\mathbf{x}^{j}_{t})$.
@@
Objects
4
5
6
7
all
Distractors
1
2
3
all
DOOM
0.998
0.975
0.974 0.958 0.976
DOOM
0.918 0.907 0.901 0.909
LOOM
0.994
0.978
0.972
0.951
0.974
LOOM
0.894
0.873
0.868
0.878
Baseline
0.938
0.786
0.765
0.702
0.798
Baseline
0.797
0.763
0.751
0.770
@@
Pose and relation prediction To get the predicted pose and relations for future steps, we utilize a latent space dynamics model to propagate the current state information to the future after multiple actions. We define the latent space dynamics to propagate the $\mathbf{x}^{i}_{t}$ to ${\mathbf{x}^{\prime}}^{i}_{t+1}$ as ${\mathbf{x}^{\prime}}^{i}_{t+1}=\delta(\mathbf{x}^{i}_{t},A_{t}(i,\theta))$. where $i$ is the ID of the object to manipulate and $\theta$ are continuous parameters parameters. We use the same learned positional embedding to encode the discrete ID, as well as an MLP $M_{\text{ac}}$ to encode $\theta$. Since we have different skills, we use a different $\delta_{l}$ and $M_{\text{ac}_{l}}$ for each skill parameter $l$. After we get ${\mathbf{x}^{\prime}}^{i}_{t+1}$, our framework decodes it to a predicted pose ${p^{\prime}}^{i}_{t+1}$ and a relation ${r^{\prime}}^{ij}_{t+1}$. Note we use ${p^{\prime}}^{i}_{t+1}$ which discriminates from $\hat{p}^{i}_{t+1}$, as ${p^{\prime}}^{i}_{t+1}$ comes from the observation $Z_{t}$ while $\hat{p}^{i}_{t+1}$ comes from the observation $Z_{t+1}$. With the ability to predict the pose and relation with latent space dynamics, even if there are occluded objects in $Z_{t+1}$, we can still estimate the pose of the occluded objects from $(Z_{t},A_{t},\delta)$. Finally, we can apply $\delta(\cdot)$ $H$ times with a sequence of actions ${A_{0},...,A_{H-1}}$ to predict states $H$ time steps ahead as ${p^{\prime}}^{i}_{t+H}$ and ${r^{\prime}}^{{ij}}_{t+H}$.
@@
Reasoning about occluded objects We explore two different approaches to reason about occluded objects. Consider object $k$ to be occluded. In our first approach, we use the predicted pose to transform the point cloud of object $k$ from time $t-1$, $O^{k}_{t-1}$, to $t$ in order to recover the missing observation as ${O^{\prime}}^{k}_{t}$. We then combine the transformed point cloud ${O^{\prime}}^{k}_{t}$ with the current observations as $Z^{\prime}_{t}={O^{1}_{t},...,{O^{\prime}}^{k}_{t},...,O^{N}_{t}}$. This allows the relational classifier to detect the relations at time $t$ even if the $k$th object disappears from the scene. We can repeat this process for any arbitrary number of occluded objects. We name this DOOM, for Direct Occluded Object Memory. We can alternatively copy the predicted latent space embedding ${\mathbf{x}^{\prime}}^{k}_{t}$ for the occluded object to the current latent state as $\mathbf{x}^{k}_{t}$ giving updated latent state $\hat{\mathbf{x}}^{\prime}_{t}=\mathbf{x}^{0}_{t},...{\mathbf{x}^{\prime}}^{k}_{t},...,\mathbf{x}^{N}_{t}$. We call this approach LOOM for Latent Occluded Object Memory. Figure 2 illustrates DOOM and LOOM.
@@
Reasoning about reappeared objects When objects become hidden, we either augment the point cloud space or the latent space with DOOM or LOOM. When occluded objects reappear, for example, the $k$th object reappears at a timestep $t$, the UVOS tracker identifies this as the same ID for the previously occluded object. We can then remove the augmented memory state associated with object $k$ from the model and instead pass this object’s observation through the point cloud encoder like normal for both DOOM and LOOM.
@@
Reasoning about novel objects’ appearance UVOS identifies novel objects when they first appear. In this case, each novel object segment receives a unique ID and is handled like any other object currently observed in the point cloud.
@@
Training details Our training loss is the sum of three terms. First, we want our model to predict the pose, relations, and environment identity at the current step $t$. Given prediction $(\hat{p}^{i}_{t},\hat{r}^{ij}_{t},\hat{e}^{i}_{t})$ and ground truth labels $(p^{i}_{t},r^{ij}_{t},e^{i}_{t})$, we get the current observation loss $L_{c}=\sum_{t=1}^{H}CE(\hat{r}^{ij}_{t},r^{ij}_{t})+||\hat{p}^{i}_{t}-p^{i}_{t}||_{2}^{2}+CE(\hat{e}^{i}_{t},e^{i}_{t})$. Second, we require our models to learn that the predicted latent space from the previous observation ${\mathbf{x}^{\prime}}^{i}_{t+x}$ should be similar to the latent space from the current time observation $\mathbf{x}^{i}_{t+x}$. We call this the latent space regularization loss, defined as $L_{ls}=\sum_{t=1}^{H}\sum_{x=t}^{H}||\mathbf{x}^{i}_{t+x}-{\mathbf{x}^{\prime}}^{i}_{t+x}||_{2}^{2}$. Last, our model should predict correct outputs from the state, ${\mathbf{x}^{\prime}}^{i}_{t+x}$ predicted by the latent dynamics; giving loss: $L_{d}=\sum_{t=1}^{H}\sum_{x=t}^{H}CE({{r^{\prime}}^{ij}_{t+x}},r^{ij}_{t+x})+||{p^{\prime}}^{i}_{t+x}-p^{i}_{t+x}||_{2}^{2}+CE({e^{\prime}}^{i}_{t+x},e^{i}_{t+x})$. We train on the full loss function $L=L_{c}+L_{ls}+L_{d}$ using the Adam optimizer.

### v3

@@ I Introduction
For robots to assist humans in their daily lives in roles such as home assistants and caregivers for elders, they must be able to reason about objects not observed in their current perceptual data. For example, if asked to retrieve an apple stored inside a cabinet, the robot should remember where it was placed and that it must first open the cabinet to grasp the object. Further if the robot picks up a box with objects in it, it should know that these objects will move with the box, while those sitting near the box will not. It is desirable for robots to maintain these models across multiple tasks in order to enable long-term autonomy.
The problem of maintaining a persistent model of a robot’s world from partial observations has a rich history in robotics, primarily in the form of mapping [1, 2, 3, 4, 5]. While some attempts have been made to incorporate object-level semantics into sensor-derived maps [6, 7, 8], these approaches primarily build immutable, monolithic models not suitable for complex manipulation planning such as multi-object rearrangement [9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. Such reasoning can enable robots to predict preconditions for actions [12], achieve language goals [13], and achieve logical goal relations [18, 15]. Thus a problem arises in how to build a consistent estimate of the world state amenable to contemporary, learning-based approaches to manipulation planning.
The standard learning-based approach to persistent memory treats the problem as a sequence prediction task, typically using autoregressive recurrent or transformer-based neural networks [19, 20, 21, 22, 23]. While these models have shown some interesting results, we hypothesize that they will have difficulty maintaining long-term memory of objects that remain occluded through many updates to a recurrent memory. One approach to overcome this issue would be to maintain the entire history of robot observations in order to use as input to a model (e.g. transformer) that could selectively attend to the relevant bits of input to determine a coherent state estimate. This is obviously untenable for long-horizon tasks as the input would grow linearly with time, inevitably becoming too large to efficiently manage.
Figure 2: Overview of our approaches. As the robot takes action over time, some objects may disappear and reappear, and some objects may newly appear in the scene. In this paper, we propose two types of object-oriented memory, called DOOM and LOOM, that enable the robot to plan with occluded and newly appeared objects. DOOM and LOOM utilize a UVOS algorithm to keep track of the current object list and update object memory slots based on the occlusion status of each object accordingly. (Best viewed in color.)
@@
In this paper we advocate for using a UVOS algorithm to explicitly manage our object-oriented memory. We hypothesize that explicit encoding of objects into the memory will be more robust than implicit autoregressive models in managing long-term history and successfully performing downstream planning. The representation should provide sufficient information to enable the prediction of inter-object and object-environment relations, while also having the ability to predict the effects of the robot’s actions on all objects in memory.
Specifically, we examine incorporating an explicit UVOS-based memory model into the framework of Huang et al. [17, 18], which was an effective framework to learn relational dynamics across varying object and environments. Key to its success is the ability to encode a variable number of objects for a given observation using a graph neural net [17] or transformer-based encoder [18]. Prediction of relations enables the model’s use in logic-based task planning [29], where relations have proved an effective means of communication between robots and humans [15, 13, 10, 30, 12, 17]. However, the existing framework assumes all relevant objects to be observable; and thus fails to successfully plan and execute in the scenarios shown in Fig. 1.
@@
To validate our proposed approaches, we test on varying tasks involving occluded objects, novel objects, and objects that reappear. We show some example tasks in Fig. 1. We show that both forms of memory-based state augmentation can reliably plan actions to change inter-object relations involving currently unobserved objects, such as placing object $A$ to the left of currently occluded object $B$. Furthermore we compare our explicit memory approach to a baseline implicit memory model in a scenario specifically designed to test the models’ long-term memory. Our experiments show that the implicit memory model fails to effectively reason about occluded objects, while our explicit, UVOS-based memory models continue to plan reliably. Our work thus represents the first successful approach of using memory models for reasoning about and planning to inter-object goal relations for an a priori unknown, variable number of objects under severe occlusion. For more real-world robot executions, additional results, and supplemental material including limitations and failure case analysis, see our website https://sites.google.com/view/rdmemory.
@@ II Related work
Reasoning about object permanence is an important capability for robot manipulation [31, 22, 32, 33]. Xu et al. [31] and Ebert et al. [22] are the first ones to propose deep learning models to reason about occluded objects and do downstream planning with the learned dynamics models. However, they assume goal images or goal configurations for planning which may not always be available from human operators. They additionally examine only planar pushing tasks with only a single moving object at each time. Our work examines a much more diverse set of tasks and skills, while also requiring only logic-based goal representations. Curtis et al. [33] propose a system with modules that can estimate affordances and properties to perform multi-step manipulation tasks with unknown and occluded objects. However, they assume complete object shape and have many engineered modules including the affordance module.
@@
Other works have also investigated using graph neural networks [42, 17, 10, 43, 44], transformers [13, 45, 46, 47], and diffusion models [9, 48, 49] to reason about and manipulate multiple objects. However, none have examined explicit memory management. Since multi-object manipulation usually requires a multi-step process, task and motion planning has shown promise in solving long-horizon, multi-object planning problem [50, 51, 52, 53, 18, 54]. These planners either require all objects to be known a priori in order to perform tracking for belief space planning [52] or ignore occlusion of unknown objects.
@@ III Approach
We assume the robot perceives the world as a point cloud, $Z_{t}$, at each timestep $t$. The robot then takes action $A_{t}$ and receives subsequent observation $Z_{t+1}$. At new observation $Z_{t+1}$, some objects may become occluded and other, new objects may appear. Based on the history of observations and actions $(Z_{0:t}$, $A_{0:t})$, we would like the robot to plan to achieve a goal, potentially involving previously observed, but currently occluded objects.
We define the goal as a logical conjunction of $M$ desired object and environment relations, $\mathbf{g}=r_{1}\land r_{2}\land...\land r_{M},r_{j}\in\mathcal{R}$, where $\mathbf{g}$ denotes the goal conjunction, $r_{j}$ represents a goal relation, and $\mathcal{R}$ denotes all possible relations. Our robot is given a set of $L$ parametric action primitives $\mathcal{A}=\{A_{1},\ldots,A_{L}\}$ where $A_{l}$ defines a skill, which has associated continuous skill parameters $\theta_{l}$. For example, a push skill is defined with parameters encoding the push direction and length, or a pick-and-dump skill is defined with parameters encoding the grasp pose and dump pose. The robot’s planning task is defined as finding skills and its parameters $\tau=((A_{0},\theta_{0}),\ldots,(A_{H-1},\theta_{H-1}))$ such that, when sequentially executed, transform the objects to satisfy all desired object and environment relations in the goal $\mathbf{g}$.
@@
Segmentation and Tracking At timestep $t$, we first perform UVOS using [28] and obtain $N$ object and environment segments $O^{i}_{t}\subset Z_{t},i=1,2,...,N$. The UVOS approach checks for new objects by examining the objectness of object proposals on key frames, when the robot finishes the execution of a skill. Then object segments that can be matched with previously tracked objects are assigned consistent IDs, while those that do not match any previous object are assigned new IDs. For non-key frames, the algorithm tracks all segments with a space-time transformer model [41] where each object has its own memory. It stores prior appearances of all objects hence can easily re-identify previously observed and disappeared objects. To familiarize the UVOS model with the robot, it is fine-tuned using annotations of the robot arm and objects from YCB-Video dataset [55]. Nonetheless, it still demonstrates generalization to novel object types, such as the tall cups (Row. 1), apple and orange (Row. 2) shown in Fig. 1 that it was never trained on. For environment segments, we use RANSAC [56] to find planar surfaces.
Pose and relation detection To jointly reason about the pose of each object and the relations between each pair of objects, we first process each segment $O^{i}_{t}$ with a point cloud encoder [57] to get a feature vector $P^{i}_{t}=E_{p}(O^{i}_{t})$. We use a learned positional embedding to encode the ID of each object as $I_{i}$. We randomly generate the ID for each object during training to improve generalization [58]. The latent space at step $t$ is denoted as $\mathbf{x}^{i}_{t}=E(P^{i}_{t}\oplus I_{i})$, where $E$ is a transformer encoding the interaction between different segments. To predict object pose and relations at the current step, we define three different decoders: the pose decoder $D_{p}$, the relation decoder $D_{r}$, and the environment identity decoder $D_{e}$. The third decoder classifies whether a given segment is movable or not (e.g. is a table or shelf) as $\hat{e}^{i}_{t}=D_{e}(\mathbf{x}^{i}_{t})$. The associated predictions of the other decoders are $\hat{p}^{i}_{t}=D_{p}(\mathbf{x}^{i}_{t})$ and $\hat{r}^{ij}_{t}=D_{r}(\mathbf{x}^{i}_{t},\mathbf{x}^{j}_{t})$.
@@
Objects 4 5 6 7 all Distractors 1 2 3 all
DOOM 0.998 0.975 0.974 0.958 0.976 DOOM 0.918 0.907 0.901 0.909
LOOM 0.994 0.978 0.972 0.951 0.974 LOOM 0.894 0.873 0.868 0.878
Baseline 0.938 0.786 0.765 0.702 0.798 Baseline 0.797 0.763 0.751 0.770
@@
Pose and relation prediction To get the predicted pose and relations for future steps, we utilize a latent space dynamics model to propagate the current state information to the future after multiple actions. We define the latent space dynamics to propagate the $\mathbf{x}^{i}_{t}$ to ${\mathbf{x}^{\prime}}^{i}_{t+1}$ as ${\mathbf{x}^{\prime}}^{i}_{t+1}=\delta(\mathbf{x}^{i}_{t},A_{t}(i,\theta))$. where $i$ is the ID of the object to manipulate and $\theta$ are continuous parameters. We use the same learned positional embedding to encode the discrete ID, as well as an MLP $M_{\text{ac}}$ to encode $\theta$. Since we have different skills, we use a different $\delta_{l}$ and $M_{\text{ac}_{l}}$ for each skill parameter $l$. After we get ${\mathbf{x}^{\prime}}^{i}_{t+1}$, our framework decodes it to a predicted pose ${p^{\prime}}^{i}_{t+1}$ and a relation ${r^{\prime}}^{ij}_{t+1}$. Note we use ${p^{\prime}}^{i}_{t+1}$ which discriminates from $\hat{p}^{i}_{t+1}$, as ${p^{\prime}}^{i}_{t+1}$ comes from the observation $Z_{t}$ while $\hat{p}^{i}_{t+1}$ comes from the observation $Z_{t+1}$. With the ability to predict the pose and relation with latent space dynamics, even if there are occluded objects in $Z_{t+1}$, we can still estimate the pose of the occluded objects from $(Z_{t},A_{t},\delta)$. Finally, we can apply $\delta(\cdot)$ $H$ times with a sequence of actions ${A_{0},...,A_{H-1}}$ to predict states $H$ time steps ahead as ${p^{\prime}}^{i}_{t+H}$ and ${r^{\prime}}^{{ij}}_{t+H}$.
@@
Reasoning about occluded objects We explore two different approaches to reason about occluded objects. Consider object $k$ to be occluded. In our first approach, we use the predicted pose to transform the point cloud of the object $k$ from time $t-1$, $O^{k}_{t-1}$, to $t$ in order to recover the missing observation as ${O^{\prime}}^{k}_{t}$. We then combine the transformed point cloud ${O^{\prime}}^{k}_{t}$ with the current observations as $(Z^{\prime}_{t}={O^{1}_{t},...,{O^{\prime}}^{k}_{t},...,O^{N}_{t}})$. This allows the relational classifier to detect the relations at $t$ even if the $k$th object is completely occluded. We can repeat this process for any arbitrary number of occluded objects. We name this DOOM for Direct Occluded Object Memory. Alternatively, we can copy the predicted latent space embedding ${\mathbf{x}^{\prime}}^{k}_{t}$ for the occluded object to the current latent state as $\mathbf{x}^{k}_{t}$ giving updated latent state $\hat{\mathbf{x}}^{\prime}_{t}=\mathbf{x}^{0}_{t},...{\mathbf{x}^{\prime}}^{k}_{t},...,\mathbf{x}^{N}_{t}$. We call this approach LOOM for Latent Occluded Object Memory. Figure 2 illustrates DOOM and LOOM.
@@
Reasoning about reappeared objects When an occluded object reappears, the UVOS tracker identifies it as the previously occluded object. We can then remove the augmented memory state associated with the object from the model and instead pass this object’s observation through the point cloud encoder like normal for both DOOM and LOOM.
@@
Reasoning about novel objects’ appearance UVOS identifies novel objects when they first appear. In this case, each novel object segment receives a unique ID and is handled the same as any other observable object.
@@
Training details Our training loss is the sum of three terms. First, we want our model to predict the pose, relations, and environment identity at the current step $t$. Given predictions $(\hat{p}^{i}_{t},\hat{r}^{ij}_{t},\hat{e}^{i}_{t})$ and ground truth labels $(p^{i}_{t},r^{ij}_{t},e^{i}_{t})$, we get the current observation loss $L_{c}=\sum_{t=1}^{H}CE(\hat{r}^{ij}_{t},r^{ij}_{t})+||\hat{p}^{i}_{t}-p^{i}_{t}||_{2}^{2}+CE(\hat{e}^{i}_{t},e^{i}_{t})$. Second, we require our models to learn that the predicted latent space from the previous observation ${\mathbf{x}^{\prime}}^{i}_{t+x}$ should be similar to the latent space from the current time observation $\mathbf{x}^{i}_{t+x}$. We call this the latent space regularization loss, defined as $L_{ls}=\sum_{t=1}^{H}\sum_{x=t}^{H}||\mathbf{x}^{i}_{t+x}-{\mathbf{x}^{\prime}}^{i}_{t+x}||_{2}^{2}$. Last, our model should predict correct outputs from the state, ${\mathbf{x}^{\prime}}^{i}_{t+x}$ predicted by the latent dynamics; giving loss: $L_{d}=\sum_{t=1}^{H}\sum_{x=t}^{H}CE({{r^{\prime}}^{ij}_{t+x}},r^{ij}_{t+x})+||{p^{\prime}}^{i}_{t+x}-p^{i}_{t+x}||_{2}^{2}+CE({e^{\prime}}^{i}_{t+x},e^{i}_{t+x})$. We train on the full loss function $L=L_{c}+L_{ls}+L_{d}$ using the Adam optimizer.